the length of rectangle is 3 cm more than the breadth. If the breadth is doubled and the length is decreased by 2 cm, then the area is increased 6 cm sq. Find the length and breadth of the original rectangle
Answer by macston(5194)

.
W=breadth; L=length=W+3
.
(W)(L)=(2W)(L-2)-6
(W)(W+3)=(2W)((W+3)-2)-6
W^2+3W=2W^2+2W-6
0=W^2-W-6
0=(W-3)(W+2)
W-3=0 OR W+2=0
W=3 OR W=-2
ANSWER 1: The original breadth was 3 cm.
.
L=W+3=3+3=6cm
ANSWER 2: The original length was 6 cm.
.
CHECK:
(W)(L)=(2W)(L-2)-6
(3cm)(6cm)=(2(3cm))(6cm-2cm)-6cm^2
18cm^2=((6cm)(4cm))-6cm^2
18cm^2=24cm^2-6cm^2
18cm^2=18cm^2